OpenStudy (anonymous):

Solve 7x2 + 28x = 49 by completing the square.

OpenStudy (anonymous):

\[0=7x^2 + 28x -49=x^2 + 4x - 7 = (x+2)^2 - 4 - 7 = (x+2)^2 - 11\] So we have \[(x+2)^2 = 11\] which gives:\[x+2 = \pm \sqrt{11}\] and this gives \[x= -2 \pm \sqrt{11}\]